OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 2x + sin x

OpenStudy (anonymous):

idk i dont do sine/cosine yet, sorry :(

OpenStudy (swissgirl):

\(cos(2x) = cos^2(x) - sin^2(x) = 2 cos^2(x) - 1 = 1 - 2 sin^2(x)\)

OpenStudy (anonymous):

thanks

OpenStudy (swissgirl):

ummm that wasnt the answer